A t.v.'s size is determined by the diagonal measure of its screen. What size t.v has dimensions 18 in by 24 in
Found 2 solutions by richwmiller, algebrahouse.com:
Answer by richwmiller(17219)

two ways
18,24
divide by 6
3,4 triangle
3,4,5 right triangle
5*6=30


d=sqrt(18^2+24^2)=30





Answer by algebrahouse.com(1659)

A right triangle is formed with legs of 18 and 24, and a hypotenuse being the diagonal.

Use the Pythagorean Theorem
a² + b² = c²
a and b are the legs
c is the hypotenuse

18² + 24² = c² {substituted 18 for a and 24 for b}
324 + 576 = c² {evaluated exponents}
900 = c² {added}
c = 30 {took the square root of each side}

The size is 30 inches

For more help from me, visit: www.algebrahouse.com
Also, you can Follow @AlgebraHouse on Twitter